John is twice Bill's age. In 5 years the sum of their age will be 34. Determine the present age for each person.
Answer by macston(5194)

.
B=Bill's age; J=Johns age=2B
.
(B+5)+(J+5)=34
B+J+10=34
B+J=24 Substitute for J
B+2B=24
3B=24
B=8
ANSWER 1: Bill is 8 years old.
.
J=2B=2(8)=16
ANSWER 2: John is 16 years old.
.
CHECK:
(B+5)+(J+5)=34
(8+5)+(16+5)=34
13+21=34
34=34